The speed of a projectile when it reaches its maximum height is 0.38 times its speed when it is at half its maximum height.What is the initial projection angle of the projectile?

1
Expert's answer
2021-11-25T16:24:08-0500

"mv_1^2\/2+mgh\/2=mgh+mv_2^2\/2"


"v_1^2+gh=2gh+v_2^2"


"v_2^2/0.38^2+gv_0^2\sin^2\alpha/(2g)=2gv_0^2\sin^2\alpha/(2g)+v_2^2"


"v_0^2\cos^2\alpha/0.38^2+gv_0^2\sin^2\alpha/(2g)=2gv_0^2\sin^2\alpha/(2g)+v_0^2\cos^2\alpha"


"\cos^2\alpha/0.38^2+\sin^2\alpha/2=\sin^2\alpha+\cos^2\alpha"


"\cos^2\alpha(1/0.38^2-1)=\sin^2\alpha-\sin^2\alpha/2"


"5.925=0.5\tan^2\alpha\to\tan\alpha=3.442\to\alpha=73.8°" . Answer